OpenStudy (anonymous):

An isosceles triangle in which angles B and C have the same measure is shown. Find the measure of each angle whose degree measure us represented with variables. Angle A: x+8y-1 Angle B: 3y+4 Angle C:7x+5

OpenStudy (anonymous):

A: 100 B:40 C:40

OpenStudy (anonymous):

7x+5=3y+4 7x-3y=-1 add all the angles...to get... 8x+11y+8=180 8x+11y=172 .......... (7x-3y=-1)*11 (8x+11y=172)*3 ........... 77x-33y=-11 24x+33y=516 .......... 101x=505 x=5 ...... 7(5)+5=40....Angle C...= Angle B Which means Angle A= 100...